# A Closed-form Solution for Weight Optimization in Fully-connected Feed-forward Neural Networks

### v1

@@ III-B Real-world Experiments
Figure 7: CA (%) versus epochs illustration, for $N=50$ with one hidden layer using MNIST database.
@@
Figure 8: CA (%) versus epochs illustration, for $N=50$ with one hidden layer using Fashion-MNIST database.
Table I summarizes the best achieved CA (%) results (on average) for the two considered datasets. The results presented in the table corroborate our claim that the proposed solution practically matches the performance of the state-of-the-art methods in both training and test phases.
@@
Dataset Algorithm BPLS in Algorithm 2 Adam NAG SGD AdaGrad
MNIST (training & test) 90.04 87.73 92.23 90.19 90.98 88.83 91.57 90.35 78.05 77.50
Fashion-MNIST (training & test) 83.99 82.57 85.97 83.77 84.07 81.98 84.81 83.59 70.69 69.94
The average running times (in seconds) are exhibited in Table II. It is worth mentioning that this analysis was performed on the machine with the following characteristics: CPU: AMD Ryzen 9 7900$\times$12-core processor $\times$ 24, RAM: $64$ GB, where the Fashion-MNIST dataset was employed and batch $=1$ was considered for the existing algorithms. The table clearly shows the superiority of the proposed BPLS algorithm over the existing ones, obtaining its results more than $1000$ times faster.
TABLE II: Average Running Time (s) per epoch of the Considered Algorithms
@@
Running Time 2.87 3578 3301 3267 3310
To conclude the results analysis, Fig. 9 illustrates the CA (%) versus average running time (s) comparison for the Fashion-MNIST test dataset, which summarizes the main findings of this section. Ideally, one would desire the results to lie in the upper-left corner of the figure. It is exactly where the results of the proposed BPLS algorithm can be found, while the results of the existing algorithms are situated in the right superior corner, illustrating the enormous difference in running times.
Figure 9: CA (%) versus average running time (s) comparison for the Fashion-MNIST test dataset.
IV Conclusions
This work proposed a novel approach to optimize weights for a fully-connected feed-forward neural network. Unlike the existing sub-optimal solutions that base their computation on iterative back-propagation and first-order gradient-based optimization schemes, the proposed solution optimizes the weights of each neuron in a single iteration and its solution is given in closed-form. It is founded on the least squares criterion, where a system of linear equations is derived by using back propagation according to the network architecture (going from right to left), where the solutions in the preceding layer are used to obtain the inputs in the succeeding one. The new algorithm is suitable for parallel implementation and can thus optimize the weights of all neurons in a single layer simultaneously. The presented numerical experiments show that the proposed solution matches the performance of the state-of-the-art methods in terms of accuracy and surpasses them significantly in terms of running time (more than $1000$ times faster in the considered network architecture).

### v2

@@ IV-B Real-world Experiments
Figure 7: CA (%) versus epochs illustration, with one hidden layer ($N_{1}=60$) using MNIST dataset.
@@
Figure 8: CA (%) versus epochs illustration, with two hidden layers ($N_{1}=100$ and $N_{2}=70$) using MNIST dataset.
[…]
(a) Training phase
(b) Test phase
Figure 9: CA (%) versus epochs illustration, with one hidden layer ($N_{1}=60$) using Fashion-MNIST dataset.
[…]
(a) Training phase
(b) Test phase
Figure 10: CA (%) versus epochs illustration, with two hidden layers ($N_{1}=100$ and $N_{2}=70$) using Fashion-MNIST dataset.
Table I summarizes the best achieved CA (%) results (on average) for the two datasets and two considered network architectures. The results presented in the table corroborate previous claim that the proposed solution practically matches the performance of the state-of-the-art methods in both training and test phases.
@@
Dataset Algorithm BPLS Adam NAG SGD AdaGrad
MNIST, $L=2$ (training & test) 89.77 89.47 94.40 92.91 94.40 92.91 90.94 90.69 57.72 58.10
MNIST, $L=3$ (training & test) 88.72 88.38 93.96 92.77 93.96 92.77 50.50 50.91 91.35 91.23
Fashion-MNIST, $L=2$ (training & test) 83.25 81.23 87.37 83.94 86.93 83.94 62.30 61.89 80.71 79.75
Fashion-MNIST, $L=3$ (training & test) 82.03 80.21 85.67 83.31 85.43 83.09 54.95 54.78 82.90 81.65
The average running times (in seconds) are exhibited in Table II. It is worth mentioning that this analysis was performed on the machine with the following characteristics: CPU: Intel Core i7-8700 CPU $3.20$GHz $\times$ $12$ with $64$GB RAM and GPU: NVIDIA GeForce RTX $2070$ SUPER/PCIe/SSE2222 where the Fashion-MNIST dataset was employed in a fully-connected network with two hidden layers, where $\text{batch}=1$ was considered for the existing algorithms. It is worth mentioning here that, by employing the existing PyTorch library, the state-of-the-art methods are fully optimized and adapted to CPU implementation. Even so, the table clearly shows the superiority of the proposed BPLS algorithm over the existing ones, obtaining its results about $20$ times faster.
TABLE II: Average Running Time (s) of the Considered Algorithms
@@
Time (CPU) 0.26 4.93 4.94 4.83 4.86
Time (GPU) 0.17 4.85 4.82 4.69 4.73
To conclude the results analysis, Fig. 11 illustrates the CA (%) versus average GPU running time (s) comparison for the Fashion-MNIST test dataset in a network with two hidden layers, which summarizes the main findings of this section. Ideally, one would desire the results to lie in the upper-left corner of the figure. It is exactly where the results of the proposed BPLS algorithm can be found, while the majority of results of the existing algorithms are situated in the right superior corner, illustrating the significant difference in running times.
Figure 11: CA (%) versus average GPU running time (s) per epoch comparison for the Fashion-MNIST test dataset in a fully-connected network with two hidden layers.
V Conclusions
This work proposed a novel approach to optimize weights for a fully-connected feed-forward neural network. Unlike the existing solutions that base their computation on iterative back-propagation and first-order gradient-based optimization schemes, the proposed solution optimizes the weights of each neuron in a single iteration and its solution is given in closed-form. It is founded on the least squares criterion, where a system of linear equations is derived by using back propagation according to the network architecture (going from right to left), where the solutions in the preceding layer are used to obtain the inputs in the succeeding one. The new algorithm is suitable for parallel implementation and can thus optimize (in the LS sense) the weights of all neurons in a single layer simultaneously. The presented numerical experiments show that the proposed solution matches the performance of the state-of-the-art methods in terms of accuracy and surpasses them significantly in terms of running time (around $20$ times faster in the considered network architecture).